Suppose the discriminating monopolist is selling a product in two separate markets in which demand function are                                                                                                                        

P1 = 80 – 2.5Q1

P2 = 180 – 10Q2

The monopolist total cost function is

TC = 50 + 40Q



1
Expert's answer
2021-05-12T15:00:27-0400

MR1 = TR'(Q1) = 80 - 5Q1,

MR2 = TR'(Q2) = 180 - 20Q2,

MC = 40,

For the first market:

80 - 5Q1 = 40,

Q1 = 8 units,

P1 = 80 - 2.5*8 = 60.

For the second market:

180 - 20Q2 = 40,

Q2 = 7 units,

P2 = 180 - 10*7 = 110.